OpenStudy (anonymous):

solve algebraically (without using logarithms): (1/8)^(2x) = 16^(x-5)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \left( 1/8\right)^{2x}=16 ^{x-5}}\) \(\color{#000000 }{ \displaystyle \left( 1/2^3\right)^{2x}=16 ^{x-5}}\) \(\color{#000000 }{ \displaystyle \left( 2^{-3}\right)^{2x}=16 ^{x-5}}\) \(\color{#000000 }{ \displaystyle \left( 2\right)^{(2x)\times (-3)}=16 ^{x-5}}\)

OpenStudy (solomonzelman):

go from there

OpenStudy (solomonzelman):

(oh, i forgot to do the same to the other side, lol)

OpenStudy (solomonzelman):

no, go ahead, if you were typing. I didn't mean to interupt

OpenStudy (solomonzelman):

\(16=2^4\). Then, the rule is\((a^b)^c=a^{b\times c}\)

OpenStudy (solomonzelman):

k...

OpenStudy (anonymous):

that wasn't me typing by the way

OpenStudy (solomonzelman):

Example problem, \(\color{#000000 }{ \displaystyle \left(\frac{1}{9} \right)^{3x+5}=81^{x} }\) We know that \(9=3^2\). \(\color{#000000 }{ \displaystyle \left(\frac{1}{3^2} \right)^{3x+5}=81^{x} }\) the rule is: \(\cfrac{1}{a^b}=a^{-b}\) \(\color{#000000 }{ \displaystyle \left(3^{-2}\right)^{3x+5}=81^{x} }\) [another] rule is: \((a^b)^c=a^{b\times c}\) \(\color{#000000 }{ \displaystyle \left(3\right)^{(3x+5)\times (-2)}=81^{x} }\) simplifying the exponent on the left, \(\color{#000000 }{ \displaystyle \left(3\right)^{-6x+10}=81^{x} }\) now, we know that \(81=3^4\) \(\color{#000000 }{ \displaystyle \left(3\right)^{-6x+10}=(3^4)^{x} }\) will once again apply: \((a^b)^c=a^{b\times c}\) \(\color{#000000 }{ \displaystyle \left(3\right)^{-6x+10}=(3)^{(x)\times (4)} }\) \(\color{#000000 }{ \displaystyle \left(3\right)^{-6x+10}=(3)^{4x} }\) and we infer that \(\color{#000000 }{ \displaystyle -6x+10=4x\quad \Rightarrow \quad 10=10x\quad \Rightarrow \quad 1=x }\)

OpenStudy (solomonzelman):

(Reply to your latest comment: I know)

OpenStudy (anonymous):

alright just finished, I understand now!

OpenStudy (solomonzelman):

Fabulous!

OpenStudy (solomonzelman):

Another example problem, \(\color{#000000 }{ \displaystyle \left(\sqrt{5}\right)^{2x+8}=\left(\sqrt{625}\right)^x }\) \(\color{#000000 }{ \displaystyle \left(5^{1/2}\right)^{2x+8}=\left(625^{1/2}\right)^x }\) \(\color{#000000 }{ \displaystyle \left(5\right)^{(2x+8)\times (1/2)}=\left(125\right)^{(x)\times (1/2)} }\) \(\color{#000000 }{ \displaystyle \left(5\right)^{x+4}=\left(625\right)^{\frac{1}{2}x} }\) \(\color{#000000 }{ \displaystyle \left(5\right)^{x+4}=\left(5^4\right)^{\frac{1}{2}x} }\) \(\color{#000000 }{ \displaystyle \left(5\right)^{x+4}=\left(5^4\right)^{\left(\frac{1}{2}x \right)\times (4)} }\) \(\color{#000000 }{ \displaystyle 5^{x+4}=5^{2x } }\) \(\color{#000000 }{ \displaystyle x+4=2x }\) \(\color{#000000 }{ \displaystyle 4=x }\)

OpenStudy (solomonzelman):

the idea here is to be able to catch any perfect Nth's when neccesary, and to know the rules of exponents.